OpenStudy (anonymous):

Hi! could someone help me solve this? (1/(1-x))+ (1/(1+x))=3 Thank you!

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[\frac{1}{x-1}+\frac{1}{x+1}=3\]??

OpenStudy (misty1212):

multiply both sides by \((x+1)(x-1)\) to clear the fractions, get \[x+1+x-1=3(x+1)(x-1)\] or just \[2x=3(x+1)(x-1)\] then solve the resulting quadratic equation for \(x\)